aaron blows across plucks a rubber band and it vibrates with a fundamental frequency of 472 hz. how long is the rubber band?
1
Expert's answer
2018-02-12T07:41:07-0500
f=v/2L
L=v/2f=330/(2(472))=0.35 m.